A bank contained $1.75 in nickels and dimes. in all, there were 26 coins in the bank. how many of each are there?
Answer by mananth(16949)

let nickels be x
dimes be y
x+y = 26............1
..
5x+10y=175..........2
multiply (1) by -5
-5x-5y=-130
add this to (2)
10y-5y=175-130
5y=45
/5
y = 9 number of dimes
x= 26-9=17 number of nickels
..
m.ananth@hotmail.ca